### v1

@@ Abstract
To tackle the crucial problem of crime, evidence at indoor crime scenes must be analyzed before it becomes contaminated or degraded. Here, as an application of artificial intelligence (AI), computer vision, and robotics, we explore how a blimp could be designed as a kind of "floating camera" to drift over and record evidence with minimal disturbance. In particular, rapid prototyping is used to develop a proof-of-concept to gain insight into what such blimps could do, manually piloted or semi-autonomously. As a result, we show the feasibility of attaching various components to an indoor blimp, and confirm our basic premise, that blimps can sense evidence without producing much wind. Some additional suggestions–regarding mapping, sensing, and path-finding–aim to stimulate the flow of ideas for further exploration.
@@
Crime is an important problem, ruining numerous lives: the global cost of violence has been estimated at a staggering $9.4 trillion USD, with homicide at $705.89 billion and assaults at $325.27 billion [1]. Furthermore, current approaches are not always successful, with low "clearance rates" the norm rather than the exception (e.g., 40% of homicides worldwide are not being solved, with worsening rates in various countries) [2]. As voiced in the UN’s 16th sustainable development goal (Peace and Justice), more needs to be done to prevent crime. However, crime often takes place behind closed doors and without witnesses, which requires careful documentation to unravel what happened. In such cases, "the crime scene investigator must set one goal above all others: secure the integrity of the crime scene . . . Lost or compromised evidence makes a crime scene investigator’s job harder and can seriously damage investigations. Blood spatter patterns and fingerprints can be inadvertently smudged, footprints or tire treads walked on if care is not taken, and trace evidence can be scattered hither and yon by those unaware of its very presence".111universalclass.com/articles/law/setting-crime-scene-perimeters.htm Evidence can also degrade naturally over time (e.g., DNA due to bacteria, heat, light, moisture, and mold) [3] when investigations are delayed due to under-staffing and under-funding–as in the saying "when seconds count, the police are only minutes away". To preserve indoor crime scenes more carefully and automatically, we imagine that blimps with cameras could be used to float over and record potentially sensitive evidence.
@@
The term "blimp" refers to a non-rigid, lighter-than-air aircraft that possesses actuators and sensors within a "gondola" (or "nacelle") to enable flight and monitoring.222https://www.airships.net/dirigible Here we use the term in a relaxed sense, permitting some small rigid structure, in the belief that our ideas could be also generally relevant for other kinds of airships, zeppelins, lighter-than-air unmanned aerial vehicles (LTA-UAVs), dirigibles, balloons, and aerostats. While aerodynes like drones could also be useful for indoor crime scene analysis, we believe that blimps feature some merits such as safety, silence, stability (ability to hover over long periods of time), sustainability (low cost, low energy/emission), soft landing, easy transport when deflated, and/or minimal training/licensing requirements for operators [4]. Conversely, the demerits of blimps, such as slow speed, large size, and sensitivity to wind, might be less problematic indoors in smaller domestic areas that are quick to cover, while still wide enough for humans to move around, and potentially less windy than large outdoor sites.
@@
A challenge is that such blimps intended to operate within difficult human environments in the context of crime would require a complex array of hardware and software capabilities. To tackle this challenge, we adopted a rapid prototyping methodology that emphasizes "flexibility, possibility, and design insights as incubated subjectively through the designer", seeking to "cast our net widely" [5]. Thus, our contribution lies in reporting on some opportunities and limitations revealed during our design experience.
@@ 2 Related Work
The basic promise of flying platforms for facilitating crime scene analysis has been explored, by manually piloting (or simulating flight of) a drone to spot mock-up bloodstains, guns, knives, and bodies [6, 7, 8]. We have also previously presented an overview of various tasks such a flying robot could perform at a crime scene [9]. However, Bucknell and Bassindale raised the alarm that downwash (wind from a drone’s propellers) could disturb sensitive evidence [10]. This could be problematic, since fibers and fine particles have been used, in some important but difficult cases, to catch various criminals, such as the "Route 40 Killer" or "Green River Killer".333https://listverse.com/2022/03/05/10-criminals-who-were-caught-thanks-to-trace-fiber-evidence/,444https://www.nbcnews.com/news/us-news/gary-ridgway-green-river-serial-killer-washington-rcna67794 In the end, Bucknell and Bassindale concluded that drones should be flown at high heights [10]. This was a consequence of how this investigation was conducted, which involved piloting a Parrot AR.Drone 2.0 at various heights. However, some wind could still be generated, and a drone might not be able to distance itself sufficiently when flying under low ceilings or detecting high-up evidence on a table or shelf. Thus, it seemed likely to us that less wind might be generated by a blimp slowly drifting or hovering overhead without using propellers, which is our focus here.
@@
More generally, the use of blimps for security applications has been considered previously: For example, blimps could be used by the police to monitor and deter crime [11].555http://www.reuters.com/article/2011/01/16/us-crime-blimp-utah-idUSTRE70F1DJ20110116 Saiki et al. proposed the use of a blimp to monitor disasters, describing control equations for a 12 m prototype with a stereo camera and LiDAR [12]. As well, the Ukrainian Armed Forces are also allegedly using tethered aerostats with infrared sensing from Aerobavovna and Kvertus that can help with surveillance and communication, act as signal relays, and launch drones when an enemy drone is detected [13]. The same article points out that China and the United States have also proposed aerostats that can launch drone swarms–an interesting idea also for crime scene analysis.
@@
Various companies have also focused on designing small indoor blimps for entertainment: For example, a patent from 1994 (US5429542A) proposes a helium-filled remote-controlled saucer toy, and various toys have been sold, such as the Megatech Blimp in the early 2000s, a fish-shaped blimp called Air Swimmers in 2011 and the NanoBlimp in 2013.666https://en.wikipedia.org/wiki/Air_Swimmer,777https://ohgizmo.com/nanoblimp-is-allegedly-worlds-smallest-rc-blimp As well, the company Festo released many imaginative creations, such as a floating manta ray called in 2007, 888https://www.youtube.com/watch?v=c3-wIICjAhE a jellyfish called the Festo AirJelly in 2008, 999https://www.youtube.com/watch?v=divLsTtA5vk a penguin called the AirPenguin in 2009, 101010https://www.youtube.com/watch?v=jPGgl5VH5go a moving origami-style blimp called the SmartInversion in 2012, 111111https://www.youtube.com/watch?v=BDiTicX8HRU a swarm of dancing spherical blimps with LEDs called the eMotionSpheres in 2014, 121212https://www.youtube.com/watch?v=5iqP1oPZ3Qw and a spherical blimp that can pick up a bottle and hand it to a person, called FreeMotionHandling, in 2016. 131313https://www.youtube.com/watch?v=dv6zQr1_C9Q
@@
Papers have also explored the design of novel kinds of indoor blimps for various applications. For example, in 2012, we built a first flying humanoid robot prototype, Angel, intended to safely interact with people; in addition to exploring how to plan safe motions and convey functional or emotional meanings through flight, we also experimented with two flight mechanisms, with wings or propellers [14]. 141414https://www.youtube.com/watch?v=4vbQ2tMOjgI/ In 2015, St-Onge et al. also made a uniquely-shaped cubic flying blimp, for indoor flight at art shows [15]. As well, in 2019, Yao et al. built a blimp to safely follow people indoors, detecting faces and two hand gestures via an RGB camera, and communicating via visual flight patterns and an LED display [16]. Furthermore, in 2019, Ferdous et al. built a blimp with low-current components, including an IR sensor and IMU, to enable usage over many hours [17]. In 2020, Seguin et al. used a convolutional neural network (CNN) to estimate the horizontal position and heading of a blimp from RGB video, along with a LiDAR to estimate height [18]. Recently, in 2024, Pham et al. described a bio-inspired blimp with fins and an onboard inertial measurement unit (IMU) [19]. Additionally, in 2024, Huang et al. designed an interesting 36" blimp with an OpenMV camera that can conduct remote measurement of important vital signs such as heart rate, respiration, or blood pressure via remote photoplethysmography; the blimp detects people with YOLO, approaches via PID control, and transmits data from stably detected foreheads [20]. Furthermore, in 2025, Xu et al. described a fish-shaped flapping-wing blimp named Cuddle-Fish that users enjoyed touching [21]. While offering insights, none of these studies considered the unique challenges of crime scene analysis by a blimp.
@@
With regard to functions for movement, documentation, and sensing, much prior work with blimps seems to have focused on control algorithms to deal with wind for particular embodiments [22]–suggesting the usefulness of checking how easy it is to control the blimp manually. As well, videos taken by a blimp’s camera would require time to watch; given that various previous work has explored the use of 3D mapping of crime scenes [23], it seemed useful if the video could be used to generate a 3D model that can be immediately inspected. Additionally, various other modalities could be used for documenting evidence, such as heat, sound, smell, or distances through solids (via radar, or WiFi signals). For example, thermal traces could reveal recent activity, like jettisoned or hidden weapons [24], or drugs.
@@
Looking further into the future, it could also be useful to check the feasibility of automatically detecting and classifying evidence, and how a blimp could plan itself how to move around to acquire data. For example, we have previously used YOLO, to detect suspicious objects [25]. Such tools could be used to gain insight into an important kind of evidence, blood. Bloodstains are fundamentally divided into three main kinds: passive drops falling due to gravity, active blood (e.g., flying from a victim’s body or from a weapon), or transfer stains (e.g., smears from a bloody hand or shoe walking through blood).151515https://www.forensicsciencesimplified.org/blood/principles.html/ While Bergman et al. recently described using a CNN to discriminate between passive drip vs. active spatter bloodstains [26], we are not aware of work that has tackled classifying between all three classes (possibly due to the sensitivity and paucity of publicly available crime data).
@@
As well, in addition to making sense of what it "sees", an autonomous blimp would also be required to itself plan how to move to acquire data. The problem of how a robot should move to cover an area is known as coverage path planning (CPP). An NP-hard problem related to the Traveling Salesman Problem, CPP is relevant to various applications, from vacuum cleaning, to lawn mowing, snow removal, search-and-rescue, and aerial or underwater terrain exploration [27]. For example, typical simplified algorithms for a vacuum cleaner robot include snaking (also known as zigzag, snake, or s-path algorithm–typically combined with a random walk), as well as random motion, and spiraling [28]. Snaking was found to be comparatively efficient for achieving high coverage faster, with less unneeded passes–at the cost of more turning and potential weakness to wheel or sensor errors that could occur in cheaper robots. In more complex scenarios where optimized algorithms are desired, the state of the art involves deep reinforcement learning (RL; e.g., for vacuum cleaners [29]). Thus, various functions could be useful, which invited exploration.
@@ 3 Methods
To start exploring potential manual and automated capabilities, a basic platform was first required. With the aim of facilitating replication, cost-effective, standard-sized, off-the-shelf components and freely available software tools were used, as described below.
@@ 3.1.1 Hardware
| Lipo Shim | 2.6 g | (15uA quiescent)222222https://www.pishop.us/product/pimoroni-zero-lipo-shim/ |
@@
Computing. To obtain and wirelessly communicate sensor data, we set up a small Raspberry Pi Zero 2 W minicomputer (6.5cm × 3.0cm) with a 1 GHz quad-core 64-bit Arm Cortex-A53 CPU, 512 MB of SDRAM, and 2.4GHz WiFi.232323https://www.raspberrypi.com/products/raspberry-pi-zero-2-w As well, to facilitate user control of the blimp, an external laptop comprising an Intel i7 CPU @ 2.60GHz with 16 GB RAM was used.
@@
Sensors. RGB Camera. A downward-facing ESP32-Cam was used to stream SVGA at 30fps from its 2 Megapixel OV2640 camera. 242424https://www.espressif.com/en/products/modules/esp32 (Various alternatives exist, such as the OpenMV camera, which although costing more, features more onboard processing capabilities.)
@@
Lidars. To detect distances in front, to the side, and below the blimp (e.g., to enable wall-following or mapping), three TF-Lunas were used. TF-Luna is a small (3.5 cm x 2.1 cm x 1.3 cm, weight 5g) Time-of-Flight (TOF) single-point LiDAR based on a 850nm Vertical Cavity Surface Emitting Laser (VCSEL), which detects distances within a range of 0.2 m to 8 m with 1 cm resolution and 6 cm accuracy.252525https://www.waveshare.com/wiki/TF-Luna_LiDAR_Range_Sensor Again, various useful alternatives existed, like the VL53l1X sensor with 4 m range, but the TF-Luna’s longer range felt like it could be useful in larger rooms or corridors.
@@
Thermal camera. Along the way, to explore possibilities for other modalities (e.g., detecting warm evidence that has recently been touched by a user), we attached a MLX90640 board. MLX90640 is a small (2.5 cm x 1.8 cm x 1.6 cm, 3.5 g) 768 pixel thermal camera with 110 x 70 field of view composed of a 24 x 32 array of IR thermal sensors that detect from -40 C to 300 C with an accuracy of 2C at a frame rate of 2 Hz.262626https://www.melexis.com/en/product/mlx90640/far-infrared-thermal-sensor-array
@@
Power. To power the microcomputer and sensors, a "LiPo Shim" was used, which can provide a maximum of 1.5 A continuous current from a 1 cell lithium polymer battery (a low battery warning occurs at 3.4 V, followed by automatic shutdown at 3.0 V).272727https://shop.pimoroni.com/products/lipo-shim With our two small batteries (1000 mAh or 700 mAh), operation was approximately 20 minutes. As well, a separate 3 V CR2 battery was used to power the actuators.
@@
Actuation and Body. To achieve motion without "reinventing the wheel", we modified an existing toy, comprising three DC motors attached to propellers and a radio receiver/transmitter. The propellers allow left and right rotation, as well as upward, downward, forward, and backward translation. To control the blimp via our own program, we connected the transmitter to our laptop via some common NPN bipolar transistors (2N2222) and an Arduino Uno.282828https://www.arduino.cc To lift the gondola, three balloons were used, each with approximately 80 cm diameter, 40 cm height, and a lift capacity of approximately 70-80 g when full (thus, 200-250 g in total). (Given the weight of helium and air (0.18 kg/m3 vs. 1.29 kg/m3), and a lift of 1.11 kg by a cubic meter of helium, this suggests there was somewhat less than a tenth of a cubic meter in each balloon.)292929https://specialtygases.messergroup.com/documents/20136/3998363/Balloon%2BHelium.pdf
During our design exploration, one observation that stood out was how tricky it can be to design small blimps with many sensors: Size. First, we found with a miniature toy unit and regular sized latex balloon that virtually no payload could be accommodated and small balloons lost helium quickly. Next, we tried using two balloons with a larger toy unit, but even though each individual sensor could be fairly light, the weight of busses, cables, tape and boards to connect and affix sensors quickly added up. Furthermore, much current was required for the microcomputer, sensors, and actuators, which in turn called for including an appropriately large battery. As well, to avoid voltage spikes from the motors resetting the microcomputer, separate power supplies were used. Thus, in the spirit of easing such restrictions and enabling freer exploration, our latest version became quite large, with three balloons (total height approximately 150 m with gondola and ballast pouch). Awkwardness. Some initial attempts to use single and double sided tape and velcro to connect the gondola and balloons didn’t provide good results, and aligning bulky components seeking to rise or fall was difficult. The gondola and balloons sometimes split, with heavy tape and velcro adhering just to one surface and ripping away material, creating unclean surfaces. Previously, working with the "Angel" blimp, we had attached hooks to a table and each balloon, but this time we wished avoid adding such weight. Instead, individual balloons were temporarily stowed under tables and desks to immobilize them. When combining balloons, we tried at one point using a low ceiling (2.3 m) in a corridor. However, this was a public space with people frequently needing to pass, and once, after returning from a quick errand, we glimpsed a passerby whacking our blimp playfully, causing the balloons to break apart. Thereafter, we propped the blimp against the bottom of a projector in our lab, which had a high ceiling (3 m+). We tied strings through each balloon’s tail section (where helium is inserted) and holes in the projector shelf, as well as to chairs below. To connect the balloons more securely, we modified some wall hooks to attach capped carbon fiber rods running up along both sides of the blimp. Even then, the top balloon sometimes flew off, so we attached set screws to both the bottom and top of the blimp to secure the rods firmly. (While we could have used a bigger balloon instead of using multiple smaller balloons, we believe this modularity is useful for exploratory design, since we can shrink or expand freely, also without wasting helium.) As well, to prevent the gondola from falling off, we made a plastic base that was firmly taped to the bottom balloon and further secured with threads going to the side hooks; then bolts and nuts were used to attach and remove the gondola. Another challenge was that the sensors had to be below the propellers, so rods were used to suspend a "second" small gondola from the first, built from light cardboard. Additionally, ballast was used to achieve neutral buoyancy in air, but we noticed that recalibration was required whenever we switched out different-sized batteries (1000 mAh or 700 mAh). As well, when testing the propellers of the gondola on a table, we were surprised when even powering motors for a brief instant was even to cause the entire gondola to spin into our laptop, causing a propeller to break; thereafter we took care to firmly tape down the gondola whenever testing it.

### v2

@@ Abstract
Crime is a critical problem—which often takes place behind closed doors, posing additional difficulties for investigators. To bring hidden truths to light, evidence at indoor crime scenes must be documented before any contamination or degradation occurs. Here, we address this challenge from the perspective of artificial intelligence (AI), computer vision, and robotics: Specifically, we explore the use of a blimp as a "floating camera" to drift over and record evidence with minimal disturbance. Adopting a rapid prototyping approach, we develop a proof-of-concept to investigate capabilities required for manual or semi-autonomous operation. Consequently, our results demonstrate the feasibility of equipping indoor blimps with various components (such as RGB and thermal cameras, LiDARs, and WiFi, with 20 minutes of battery life). Moreover, we confirm the core premise: that such blimps can be used to observe crime scene evidence while generating little airflow. We conclude by proposing some ideas related to detection (e.g., of bloodstains), mapping, and path planning, with the aim of stimulating further discussion and exploration.
@@
Crime remains a global crisis, ruining numerous lives and draining public resources: Intentional homicides cruelly took approximately half of a million lives in 2021—the highest rate in twenty years [1]. The global cost of violence has also been estimated at a staggering $9.4 trillion USD, with homicide at $705.89 billion and assaults at $325.27 billion [2]. Furthermore, current countermeasures are often unsuccessful, with 40% of homicides worldwide going unsolved, and worsening clearance rates noted in countries such as the United States (US), Canada, Trinidad, and Tobago [3]. As emphasized in the United Nation’s 16th sustainable development goal (Peace and Justice), new approaches are urgently needed to improve crime prevention and investigation [1].
@@
Yet, crime often takes place indoors—behind closed doors and without witnesses—requiring careful, often painstaking documentation to piece together what transpired. In such cases, as noted in forensic literature, "the crime scene investigator must set one goal above all others: secure the integrity of the crime scene . . . Lost or compromised evidence makes a crime scene investigator’s job harder and can seriously damage investigations. Blood spatter patterns and fingerprints can be inadvertently smudged, footprints or tire treads walked on if care is not taken, and trace evidence can be scattered hither and yon by those unaware of its very presence".111https://universalclass.com/articles/law/setting-crime-scene-perimeters.htm Furthermore, time is a critical factor: Evidence can degrade due to bacteria, heat, light, moisture, or mold [4], and investigations are often delayed or cut short by understaffing or limited resources. As the saying goes, "when seconds count, the police are only minutes away." Thus, to help preserve indoor crime scenes more carefully and automatically, we imagine that blimps with cameras could be used to float over and record potentially sensitive evidence.
@@
The term "blimp" traditionally refers to a non-rigid, lighter-than-air aircraft equipped with actuators and sensors within a “gondola” or “nacelle” to enable flight and monitoring.222https://www.airships.net/dirigible Here, we adopt the term in a relaxed sense—e.g., allowing for some small rigid structure—in the belief that our ideas could be also broadly relevant for other kinds of aerostats, airships, dirigibles, and lighter-than-air unmanned aerial vehicles (LTA-UAVs), like hybrid dynastats or rotastats, zeppelins, and balloons. While aerodynes such as drones could also aid in indoor crime scene analysis, blimps offer some unique advantages: safety, silence, stability (ability to hover over extended periods), sustainability (due to low cost, energy use, and emissions), soft landing, easy transport when deflated, and minimal training or licensing requirements for operators [5]. Conversely, typical disadvantages—such as sensitivity to wind, low speed, and large size—could be less limiting in indoor environments, which are generally less windy and more compact, yet wide enough to be navigable by humans.
@@
A challenge is that blimps intended to help solve difficult crimes in chaotic human environments would require a complex array of hardware and software capabilities. To tackle this challenge, we adopted a rapid prototyping methodology that emphasizes "flexibility, possibility, and design insights as incubated subjectively through the designer", seeking to "cast our net widely" [6]. Thus, our contribution lies in reporting on some opportunities and limitations revealed during our design experience.
@@ 2 Related Work
The basic potential of flying platforms for facilitating crime scene analysis has been explored, by manually piloting drones—or simulating their flight—to spot mock-up bloodstains, guns, knives, and bodies [7, 8, 9]. We have also previously presented an overview of tasks that a flying robot could perform at a crime scene [10]. However, Bucknell and Bassindale raised the alarm that downwash—wind from a drone’s propellers—could disturb sensitive evidence [11]. This could be problematic, since fibers and fine particles have played a crucial role in solving some important but challenging cases, such as those of the "Route 40" or "Green River" killers.333https://listverse.com/2022/03/05/10-criminals-who-were-caught-thanks-to-trace-fiber-evidence/,444https://www.nbcnews.com/news/us-news/gary-ridgway-green-river-serial-killer-washington-rcna67794 In their study, Bucknell and Bassindale concluded that drones should be flown at higher altitudes, based on experiments involving a Parrot AR.Drone 2.0 at varying heights [11]. Nonetheless, wind could still be generated, especially in indoor settings with low ceilings or when examining elevated evidence on tables or shelves. In contrast, a blimp that hovers or drifts slowly without using its propellers could generate less airflow—an assumption that motivates the present study.
@@
More broadly, blimps have been previously considered for security-related applications: For instance, Murphy and colleagues described how the police could use blimps to monitor and deter crime—an idea discussed in the city of Ogden in the US [12].555https://www.blimpinfo.com/uncategorized/plans-for-ogden-police-blimp-canceled/ Saiki et al. also proposed the use of a blimp to monitor disasters, describing control equations for a 12 m prototype with a stereo camera and LiDAR [13]. As well, the Ukrainian Armed Forces are using tethered aerostats with infrared sensing from Aerobavovna and Kvertus to help with surveillance, communication, and drone deployment when an enemy drone is detected [14]. The same article notes that China and the US have also proposed aerostats that can launch drone swarms—an idea that could also hold promise for analysis of larger crime scenes.
@@
Various companies have also designed small indoor blimps, for entertainment: For example, a 1994 patent (US5429542A) proposed a helium-filled, remote-controlled saucer toy; over the years, various toys have followed, such as the Megatech Blimp in the early 2000s, a fish-shaped blimp called Air Swimmers in 2011 and the NanoBlimp in 2013.666https://en.wikipedia.org/wiki/Air_Swimmer,777https://ohgizmo.com/nanoblimp-is-allegedly-worlds-smallest-rc-blimp As well, the company Festo has stirred interest with many imaginative creations: e.g., Air_ray, a floating manta ray introduced in 2007;888https://www.youtube.com/watch?v=c3-wIICjAhE AirJelly, a jellyfish from 2008;999https://www.youtube.com/watch?v=divLsTtA5vk AirPenguin, a penguin from 2009;101010https://www.youtube.com/watch?v=jPGgl5VH5go SmartInversion, a moving origami-style blimp from 2012;111111https://www.youtube.com/watch?v=BDiTicX8HRU eMotionSpheres, a swarm of dancing spherical blimps with LEDs from 2014;121212https://www.youtube.com/watch?v=5iqP1oPZ3Qw and FreeMotionHandling, a spherical blimp that can pick up a bottle and hand it to a person, from 2016.131313https://www.youtube.com/watch?v=dv6zQr1_C9Q
@@
Several research papers have also explored the design of novel indoor blimps for diverse applications: For example, in 2012, we built a first flying humanoid robot prototype, Angel, intended to safely interact with people [15].141414https://www.youtube.com/watch?v=4vbQ2tMOjgI/ In addition to exploring how to plan safe motions and convey functional or emotional meanings through flight, we also experimented with two flight mechanisms, utilizing wings or propellers. In 2015, St-Onge et al. made a uniquely-shaped cubic flying blimp, for indoor flight at art shows [16]. As well, in 2019, Yao et al. built a blimp to safely follow people indoors, detecting faces and two hand gestures via an RGB camera, and communicating via visual flight patterns and an LED display [17]. That same year, Ferdous et al. built a blimp with low-current components, including an infrared (IR) sensor and inertial measurement unit (IMU), to enable usage over many hours [18]. In 2020, Seguin et al. used a convolutional neural network (CNN) to estimate the horizontal position and heading of a blimp from RGB video, along with a LiDAR to estimate height [19]. Recently, in 2024, Pham et al. described a bio-inspired blimp with fins and an onboard IMU [20]. Additionally, in 2024, Huang et al. designed an interesting 36" blimp with an OpenMV camera that can conduct remote measurement of important vital signs such as heart rate, respiration, or blood pressure via remote photoplethysmography; the blimp detects people with YOLO, approaches via PID control, and transmits data from stably detected foreheads [21]. Furthermore, in 2025, Xu et al. described a fish-shaped, flapping-wing blimp named Cuddle-Fish that users enjoyed touching [22]. The same year, Hong and Tanaka presented unique walking humanoids with balloon torsos and articulated legs containing LEDs and various sensors, that can be controlled with gamepads or directed via external fans at an art installation [23]. While informative, none of these studies considered the unique challenges of crime scene analysis by a blimp.
@@
With regard to functions for movement, documentation, and sensing to enable manual piloting, much prior work on blimps seems to have focused on control algorithms to deal with wind for particular embodiments [24]—suggesting the usefulness of checking how easy it is to pilot a CSA blimp manually. As well, videos captured by a blimp’s camera require time to watch; given that various previous work has explored the 3D mapping of crime scenes [25], it seemed useful if a video could be leveraged to generate a 3D model that can be immediately inspected. Additionally, various other modalities could be used for documenting evidence, such as heat, sound, smell, or distances through solids (via radar or WiFi signals). For example, thermal traces could reveal recent activity, like jettisoned or hidden weapons [26] or drugs.
@@
Looking further ahead, it could also be beneficial to explore the feasibility of automatically detecting and classifying evidence, and how a blimp itself could plan to move to acquire data. For example, we have previously used YOLO, to detect suspicious objects [27]. Such a tool could also be used to gain insight into an important kind of evidence, blood. Bloodstains are generally categorized into three main kinds: passive (e.g., from drops falling under gravity), active (e.g., from blood expelled from a victim’s body or a weapon), or transfer (e.g., from contact by a bloody hand or shoe).151515https://www.forensicsciencesimplified.org/blood/principles.html While Bergman et al. recently described using a CNN to discriminate between passive drip vs. active spatter bloodstains [28], we are not aware of work that has tackled classification across all three categories—possibly due to the sensitivity and scarcity of publicly available crime data.
@@
In addition to interpreting what it "sees", an autonomous blimp should also plan how to move to acquire data. The problem of how a robot should move to cover a given area is known as coverage path planning (CPP). An NP-hard problem related to the Traveling Salesman Problem, CPP arises in a wide range of applications, from vacuum cleaning, to lawn mowing, snow removal, search-and-rescue, and aerial or underwater terrain exploration [29]. For example, typical simplified algorithms used by robotic vacuum cleaners include snaking (also known as a zigzag or s-path algorithm—often combined with random walks), as well as spiraling and random motion [30]. Among these, snaking was found to be efficient at achieving high coverage faster, with fewer unneeded passes—at the cost of increased turning and weakness to wheel or sensor errors that can occur in lower-cost robots. In more complex scenarios where optimized algorithms are desired, the state of the art increasingly involves deep reinforcement learning (RL; e.g., for vacuum cleaners [31]). Accordingly, various manual and autonomous capabilities appeared like they could be useful, inviting exploration.
@@ 3 Methods
To start exploring potential capabilities, a basic platform was first required. With the aim of facilitating replication, cost-effective, standard-sized, off-the-shelf components and freely available software tools were used, as described below.
@@ 3.1.1 Hardware
| Lipo Shim | 2.6 g | (15 A quiescent)222222https://www.pishop.us/product/pimoroni-zero-lipo-shim/ |
@@
Computing. To obtain and wirelessly communicate sensor data, we set up a small Raspberry Pi Zero 2 W minicomputer (6.5 cm × 3.0 cm) with a 1 GHz quad-core 64-bit Arm Cortex-A53 CPU, 512 MB of SDRAM, and 2.4 GHz WiFi.232323https://www.raspberrypi.com/products/raspberry-pi-zero-2-w As well, to control the blimp, an external laptop comprising an Intel i7 CPU @ 2.60GHz with 16 GB RAM and a 6GB GPU was used.
@@
Sensors. RGB Camera. A downward-facing ESP32-Cam was used to stream SVGA at 30fps from its 2 Megapixel OV2640 camera.242424https://www.espressif.com/en/products/modules/esp32 (Various alternatives exist, such as the OpenMV camera,252525https://openmv.io which although costing more, features more onboard processing capabilities.)
@@
LiDARs. To detect distances in front, to the side, and below the blimp (e.g., allowing wall-following or mapping), three TF-Lunas were used. TF-Luna is a small (3.5 cm x 2.1 cm x 1.3 cm, weight 5g) Time-of-Flight (TOF) single-point LiDAR based on an 850 nm Vertical Cavity Surface Emitting Laser (VCSEL), which detects distances within a range of 0.2 m to 8 m with 1 cm resolution and 6 cm accuracy.262626https://www.waveshare.com/wiki/TF-Luna_LiDAR_Range_Sensor Again, alternatives existed, like the VL53l1X sensor with 4 m range—yet the TF-Luna’s longer range felt like it could be useful in larger rooms or corridors.
@@
Thermal camera. Along the way, to explore possibilities for other modalities (e.g., detecting warm evidence that has recently been touched by a user), we attached an MLX90640 board. MLX90640 is a small (2.5 cm x 1.8 cm x 1.6 cm, 3.5 g) 768 pixel thermal camera with a 110 x 70 field of view, composed of a 24 x 32 array of IR thermal sensors that detect from -40C to 300C with an accuracy of 2C at a frame rate of 2 Hz.272727https://www.melexis.com/en/product/mlx90640/far-infrared-thermal-sensor-array More powerful but costly alternatives such as FLIR exist.
@@
Power. To power the microcomputer and sensors, a "LiPo Shim" was used, which provides a maximum of 1.5 A continuous current from a 1 cell lithium polymer battery (a low battery warning occurs at 3.4 V, followed by automatic shutdown at 3.0 V).282828https://shop.pimoroni.com/products/lipo-shim With our two small batteries (700 mAh or 1000 mAh, approx. 15-20 g), operation was approximately 20 minutes. As well, a separate 3 V CR2 battery was used to power the actuators.
@@
Actuation and Body. To achieve motion without "reinventing the wheel", we modified an existing toy, comprising three DC motors attached to propellers and a radio receiver/transmitter. The propellers allow left and right rotation, as well as upward, downward, forward, and backward translation. To control the blimp via our own program, we connected the transmitter to our laptop via some common NPN bipolar transistors (2N2222) and an Arduino Uno.292929https://www.arduino.cc To lift the gondola, three balloons were used, each with approximately 80 cm diameter, 40 cm height, and a lift capacity of approximately 70-80 g when full (thus, 200-250 g in total). (Given the weight of helium and air (0.18 kg/m3 vs. 1.29 kg/m3), and a lift of 1.11 kg by a cubic meter of helium, this suggests there was somewhat less than a tenth of a cubic meter in each balloon.)303030https://specialtygases.messergroup.com/balloon-helium
During our design exploration, we observed that equipping small blimps with many sensors posed challenges: Size. Miniature toy units with regular latex balloons offered minimal payload capacity and lost helium quickly. Using two balloons with a larger toy improved lift, but the cumulative weight of sensors, boards, cables, tape, and power systems added up fast. High current demands from the microcomputer, sensors, and actuators required a large battery, and separate power supplies were adopted to avoid voltage spikes that could reset the microcontroller. To overcome these constraints and enable free exploration, our latest version became substantially larger, using three balloons (total height 150 cm with gondola and ballast).
Assembly. Initial attachment methods using tape or velcro proved unreliable; gondolas detached or tore balloon surfaces. Unlike our earlier “Angel” design, where balloons were anchored to a table, we avoided heavy hooks this time by temporarily tucking balloons under tables to immobilize them. At one point, we used a 2.3 m corridor ceiling to assemble the blimp, but public access made it impractical—once, we glimpsed a passerby whacking our blimp playfully, causing the balloons to break apart. We later secured the blimp under a ceiling-mounted projector in our lab, tying balloon tails also to chairs as needed, and connecting balloons via capped carbon fiber rods held by modified wall hooks. Despite this, the top balloon sometimes detached, so we added set screws at both ends to secure the rods.
To prevent the gondola from detaching, we taped a plastic base to the bottom balloon, reinforced with threads to side hooks, and affixed the gondola using bolts and nuts. Sensors had to be mounted below the propellers, so we suspended a second lightweight cardboard gondola from the main one using rods. We also used ballast to fine-tune buoyancy, but noticed that battery changes (e.g., between 700 mAh and 1000 mAh) required recalibration. While testing propellers on a tabletop, we were also surprised when only briefly powering a motor caused the entire gondola to spin and hit a laptop, breaking a propeller—after which we made sure to tape down the gondola securely when testing.